1. Hahn Signage Co. produces customized billboards and neon signs. The company uses a job costing system for costing its products. The following information is available for the month of March Cost incurred Job Added in March Direct materials Direct labour March1 Status at March 31 NS-Astarspa $25,320 SS-SamCafe BB-BobHotel BB-Alta.vac $2,500 $2,952 Completed and sold 850 10,500 3,240Completed but not sold 3,348Not completed 12,500 500 900 10440 Started and not completed Direct materials inventory Balance, March1 March purchases Balance, March 31 $ 4,250 25,300 5,700 Finished goods inventory Beginning inventory job BB-Cal-Stampede $22,100 Hahn uses a normal costing system and applies manufacturing overhead at a rate of $75 per direct labour hour. The average direct labour wage is $18 per hour. What is the total cost of job SS-SamCafe on the job cost sheet on March 31? a) $12,440 b) $15,340 c) $25,940 d) $28,840
Answer #1

Answer

  • Total Cost of Job SS Sam café = Beginning WIP Inventory balance + Direct Material + Direct Labor + Manufacturing Overhead.
  • Note:
    SS SamCafe Direct Labor cost =$ 3240
    Labor rate = $ 18 per hour
    No. of labor hours = 3240 / 18 = 180 labor hours
  • Direct Material Inventory used for Job SS SamCafe calculation:

Total material used = beginning + Purchases – Ending ba;ance
= 4250 + 25300 – 5700
= $ 23,850 = total added in March

Direct material for SS SamCafe = 23850 – 2500 – 12500 – 500 = $ 8350

  • Total cost for Job SS Sam Café = Beginning WIP Inventory balance + Direct Material + Direct Labor + Manufacturing Overhead.
    $ 850 + $ 8350 + $ 3240 + (180 hours x $ 75)
    = 850 + 8350 + 3240 + 13500
    = $ 25,940
  • Correct Answer = $ 25,940 = Option #3